OpenStudy (anonymous):

x/-6=-12 (Points : 1) No solution –2 or 2 –72 or 72 72

OpenStudy (anonymous):

x needs to be isolated to do that you need to do the reciprocal of division (multiplication) -6×-12=? *2 negatives will make a positive)

OpenStudy (jhannybean):

\[\sf \frac{x}{-6}=\frac{-12}{1}\]Cross multiply.\[\sf (1)(x) = (-6)(-12)\]

OpenStudy (anonymous):

is it negative or positive 72

OpenStudy (jhannybean):

When you mutiply two -ves you get a +ve.

OpenStudy (anonymous):

it'll only be one of them 2 negatives will make a positive

OpenStudy (jhannybean):

\[-(-72) = ~?\]